OpenStudy (anonymous):

solve for x, step by step with factoring no quad formula plz x^4 - 5x^2 - 6 = 0

OpenStudy (amistre64):

say: x^2 = a, and solve for a using quadratic methods

OpenStudy (anonymous):

@jim_thompson5910 help here :D

jimthompson5910 (jim_thompson5910):

exactly, but I'd use z instead of 'a' ('a' is its own word pretty much, which leads to confusion) to let z = x^2 So z^2 = x^4 which means x^4 - 5x^2 - 6 = 0 becomes z^2 - 5z - 6 = 0

OpenStudy (anonymous):

if i were do to this it would be (x^2-6) (x^2+1)=0

jimthompson5910 (jim_thompson5910):

(x^2-6) (x^2+1)=0 is the correct factorization

OpenStudy (anonymous):

ok then why is it that (x^2-6) cant be factored further like (x=3) (x-3)

jimthompson5910 (jim_thompson5910):

because (x+3) (x-3) = x^2 - 9

OpenStudy (anonymous):

i mean (x+3) (x-2)

jimthompson5910 (jim_thompson5910):

(x+3) (x-2) = x^2 + x - 6

OpenStudy (anonymous):

so then how do i know when to take the square root

jimthompson5910 (jim_thompson5910):

x^2-6 = 0 x^2 = 6 x = +-sqrt(6)

jimthompson5910 (jim_thompson5910):

that's one piece to it

OpenStudy (anonymous):

my problem was that after i reached the part where i factored the (x^2-6), i dont know whether to factor it or take out its square root

OpenStudy (anonymous):

still dont get it if you can explain ill appreciate it

jimthompson5910 (jim_thompson5910):

(x^2-6) (x^2+1)=0 x^2-6=0 or x^2+1=0 Now solve each equation separately

jimthompson5910 (jim_thompson5910):

you can factor x^2 - 6, but it will involve square roots, so why not just jump straight to the answer and solve for x